# BounTCHA: A CAPTCHA Utilizing Boundary Identification in AI-extended Videos

### v1

@@ Abstract.
In recent years, the rapid development of artificial intelligence (AI) especially multi-modal Large Language Models (MLLMs), has enabled it to understand text, images, videos, and other multimedia data, allowing AI systems to execute various tasks based on human-provided prompts. However, AI-powered bots have increasingly been able to bypass most existing CAPTCHA systems, posing significant security threats to web applications. This makes the design of new CAPTCHA mechanisms an urgent priority. We observe that humans are highly sensitive to shifts and abrupt changes in videos, while current AI systems still struggle to comprehend and respond to such situations effectively. Based on this observation, we design and implement BounTCHA, a CAPTCHA mechanism that leverages human perception of boundaries in video transitions and disruptions. By utilizing AI’s capability to expand original videos with prompts, we introduce unexpected twists and changes to create a pipeline for generating short videos for CAPTCHA purposes. We develop a prototype and conduct experiments to collect data on humans’ time biases in boundary identification. This data serves as a basis for distinguishing between human users and bots. Additionally, we perform a detailed security analysis of BounTCHA, demonstrating its resilience against various types of attacks. We hope that BounTCHA will act as a robust defense, safeguarding millions of web applications in the AI-driven era.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Authenticationccs: Security and privacy Access controlccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Computer vision
@@ 1. Introduction
The primary purpose of CAPTCHAs is to prevent web crawlers and bots (Kubicek et al., 2024) that simulate human behavior from sending frequent requests, which can make the operational data of web services more authentic, resistant to Sybil attacks (Douceur, 2002), and place significant strain on a server’s network. A large volume of such requests may be perceived as a denial-of-service (DoS) attack (Jhaveri et al., 2012), and attackers with multiple nodes may launch a distributed denial-of-service (DDoS) attack (Gasti et al., 2013; Karami et al., 2016). This disrupts the normal use of web applications for legitimate human users. Additionally, bots can attack social media content by promoting keywords to get them trending on X (previously Twitter), thus creating false trends that can reach a broad audience (Elmas, 2023; Hays et al., 2023). The situations above are not only happened in the centralized web, but also and more severe in the decentralized web (Zhang and Lee, 2019; Li et al., 2009). While there are various methods to detect (Zhang et al., 2022; Kang et al., 2010; Senol et al., 2024) and mitigate (Du et al., 2024; Brustoloni, 2002; Huang et al., 2023) such attacks, CAPTCHAs remain one of the most cost-effective and widely used defenses.
@@
With the recent surge in multi-modal Large Language Models (MLLMs) (Iong et al., 2024; Yin et al., 2023), artificial intelligence (AI) has gained unprecedented capabilities in handling a wide range of complex tasks, including but not limited to understanding text, multimedia, and generating synthesized conclusions. Furthermore, by granting AI the ability to control computers (Ma et al., 2024), these systems can now take autonomous actions in the digital world. An AI-powered agent bot could operate a computer much like a human, posing a significant security threat to existing web applications. Once an AI bot learns how to overcome CAPTCHA challenges, the current CAPTCHA systems could become obsolete overnight.
@@
In recent years, AI-driven video generation models have emerged in an exponential manner, with examples such as Stable Video Diffusion (SVD) (Blattmann et al., 2023), Sora [^1] https://openai.com/index/sora/ (Liu et al., 2024), Pika Labs [^2] https://pika.art/, Runway [^3] https://runwayml.com/ (Cui et al., 2024), Stability AI [^4] https://stability.ai/, Kling [^5] https://kling.kuaishou.com/, and others (Ho et al., 2022a, b; Luo et al., 2023; Khachatryan et al., 2023). The videos generated by these models have proliferated across various social media platforms. As model architectures continue to improve and the size of model parameters expands, the length and quality of the generated videos are expected to reach new heights, making it increasingly difficult to distinguish them from real content. In addition to generating videos from prompts and related images, some models are especially capable of video extension (known as video prolongation, and video prediction). Moreover, these models can customize extended scenes, actions, visuals, and objects within the video using text prompts. They can also create visuals that do not exist in the real world, similar to special effects in films. And in the research (Hénaff et al., 2019), it helps explain changes in videos that violate the predictive perception are easily noticed by humans, highlighting how humans detect unnature changes in the video content. As a result, humans are able to identify the part of a video that has been AI-extended by observing strange changes in the video.
@@
Motivation. In light of the above observations, we argue that the human ability to recognize sudden events, transitions, or anomalies inconsistent with the real world in video content can be harnessed to design a novel web CAPTCHA system. This system would serve as a defense against bot attacks on the internet. The video data used for the CAPTCHA could be generated using AI techniques, specifically by leveraging the extending video capabilities of video generation models.
@@
Approach. We design a novel CAPTCHA mechanism aimed at defending against current and future, more intelligent web robots and crawlers. This mechanism is based on abrupt changes and transitions in video frames and storylines, where users are asked to identify the points of transition, allowing to distinguish between human users and robots. Since collecting large datasets of real videos with significant transitions is costly, and the degree of transition in real videos is limited, we use AI video generation models to extend these transitions. As a result, our CAPTCHA mechanism relies on human perception to identify AI-extended video boundaries, and which we name BounTCHA. Based on this design, we develop a prototype whereby, after watching the video, users are asked to drag the progress bar to the point where they believe the transition occurs and submit their response. The collected timing data is used to determine the effective range within which humans can recognize the boundary, which then serves as the basis for distinguishing human users from robots. Following this, we conduct a detailed security analysis to assess the effectiveness of this defense method.
@@
The following sections review relevant literature, outline video data preparation, detail the prototype and studies on human performance, discuss the in-depth security analysis and conclude with limitations and future work.
@@ 2.3. Video Understanding
In the realm of video understanding tasks, video foundation models (VFMs) have embarked on promising endeavors in exploring model architectures, as evidenced by references such as (Wang et al., 2023b; Lei et al., 2021; Xu et al., 2024). Additionally, they have made significant progress in learning paradigms, (Li et al., 2023; Xu et al., 2021; Zellers et al., 2021; Xiao et al., 2024; Yao et al., 2024) serving as examples. As for the latest advancements in video understanding, Tarsier (Wang et al., 2024) emerges as a prominent family of large-scale video-language models. These models are meticulously designed to craft high-quality video descriptions. Notably, the model, along with its code and data, has been made publicly accessible for inference, evaluation, and deployment purposes. As of September 3rd, 2024, extensive evaluation results have demonstrated the remarkable capabilities of Tarsier in general video understanding. It has achieved state-of-the-art (SOTA) performance across six open benchmarks, solidifying its position as a leading approach to understand videos. Consequently, in our current research endeavor, we have decided to utilize this model to participate in the prompt generation of AI-extended videos.
@@ 3. Video Data Preparation (RQ1)
3.1. Video Generated Pipeline
@@ 3.2.2. Prompts for Video Generation Prompts
After obtaining the description and five keywords of the raw video, we incorporate this information into the subsequent video generation prompts, along with the last frame of the raw video. These inputs are then fed into the MLLM to generate a completely new video narrative. This process ensures that the generated prompts significantly differ from the storyline and visuals of the raw video, while still being grounded in its foundational elements. Subsequently, the prompts used for video generation are input into the video generative AI to contribute to the creation of the AI-extended video.
@@ 3.3. Video Quality & Shift Cutting
3.4. Video Production Time Consumption
Figure 5 illustrates the time consumption in various stages of video preparation, such as video understanding for 21.7 seconds, prompt generation for 5.4 seconds, video generation for 521.2 seconds and video compression for 3.1 seconds.
Figure 5. The time cost of the video preparation pipeline. The length of the blocks in the figure is not drawn to scale based on the time duration.
@@ 5.1. Experiment
Ethics. The school’s Institutional Review Board (IRB) reviewed and approved this human-subjects research. In order to preserve the double-blind review process, the approval number will be provided in the camera-ready version.
Participants. We recruited 186 participants for the experiment by posting flyers on campus and advertising on social media (comprising 83 participants in-person and 103 participants via ZOOM for online experiments). Participants’ ages ranged from 18 to 48 years ($\mu$ = 26.93, $\sigma$ = 7.57), and all had prior experience with web-based CAPTCHAs. Each participant compensated with \$0.7 USD.
@@ 5.2. Results
Figure 11. Left: the time bias range and the success rate of the group in the age 18-30. Mid: in the age 30-40. Right: in the age 40-50.
@@ 6.3. Multi-modal LLM (MLLM) Attack
7. Limitations and Future Work
In this section, we point out some limitations of this work, and identify future research directions.
Participants’ diversity. Since the participants we recruited are from a university setting, they are either currently pursuing or have already completed higher education, which may result in a lack of diversity among the participants. Future research could expand the range of participants’ educational levels to better assess the generalizability of this type of CAPTCHA.
Bidirectional extension. The mainstream direction of video extension is forward-only, but Sora supports bidirectional extensions [^8] https://openai.com/index/video-generation-models-as-world-simulators/, which include both forward and backward. Nevertheless, Sora is currently not available to the public. Research can be conducted on bidirectional extended videos without telling users which part is generated. We can also design the derived CAPTCHA mechanisms from it.
@@ 8. Conclusion
We designed and implemented a novel CAPTCHA, named BounTCHA, to address the growing threat posed by increasingly intelligent AI-powered bots. BounTCHA was designed to differentiate between genuine users and bots based on human perception and identification of time boundaries in video transitions and abrupt changes. We developed a prototype of BounTCHA to conduct experiments and determine the effective range of human perceptual time bias, which serves as a basis for distinguishing between real users and bots. A comprehensive security analysis was then performed on BounTCHA, covering random attacks, database attacks, and multi-modal LLM attacks. The results of the analysis demonstrated that BounTCHA effectively defends against various attack vectors. We envision BounTCHA as a robust shield in web security, safeguarding millions of web applications from AI-powered bot threats in an era where machines are becoming increasingly intelligent.

### v2

# BounTCHA: A CAPTCHA Utilizing Boundary Identification in Guided AI-extended Videos

@@ Abstract.
In recent years, the rapid development of artificial intelligence (AI) especially multi-modal Large Language Models (MLLMs), has enabled it to understand text, images, videos, and other multimedia data, allowing AI systems to execute various tasks based on human-provided prompts. However, AI-powered bots have increasingly been able to bypass most existing CAPTCHA systems, posing significant security threats to web applications. This makes the design of new CAPTCHA mechanisms an urgent priority. We observe that humans are highly sensitive to shifts and abrupt changes in videos, while current AI systems still struggle to comprehend and respond to such situations effectively. Based on this observation, we design and implement BounTCHA, a CAPTCHA mechanism that leverages human perception of boundaries in video transitions and disruptions. By utilizing AI’s capability to expand original videos with prompts, we introduce unexpected twists and changes to create a pipeline for generating guided short videos for CAPTCHA purposes. We develop a prototype and conduct experiments to collect data on humans’ time biases in boundary identification. This data serves as a basis for distinguishing between human users and bots. Additionally, we perform a detailed security analysis of BounTCHA, demonstrating its resilience against various types of attacks. We hope that BounTCHA will act as a robust defense, safeguarding millions of web applications in the AI-driven era.
@@
ccs: Security and privacy Authenticationccs: Security and privacy Access controlccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Computer vision
@@ 1. Introduction
The primary purpose of CAPTCHAs is to prevent web crawlers and bots (Kubicek et al., 2024) that simulate human behavior from sending frequent requests, which can make the operational data of web services more authentic, resistant to Sybil attacks (Douceur, 2002), and place significant strain on a server’s network. A large volume of such requests may be perceived as a denial-of-service (DoS) attack (Jhaveri et al., 2012), and attackers with multiple nodes may launch a distributed denial-of-service (DDoS) attack (Gasti et al., 2013; Karami et al., 2016). This disrupts the normal use of web applications for legitimate human users. A study on web bots found that, on average, each website used for the experiment receives over 37K requests per month and more than 50% of them attempting to bruteforce attacks and exploit vulnerabilities (Li et al., 2021). Additionally, bots can attack social media content by promoting keywords to get them trending on X (previously Twitter), thus creating false trends that can reach a broad audience (Elmas, 2023; Hays et al., 2023). It is also possible to disrupt the content recommendation systems of social media platforms in order to influence public opinion (Chen et al., 2024). The situations above are not only happened in the centralized web, but also and more severe in the decentralized web (Zhang and Lee, 2019; Li et al., 2009). While there are various methods to detect (Zhang et al., 2022; Kang et al., 2010; Senol et al., 2024; Li et al., 2023a) and mitigate (Du et al., 2024; Brustoloni, 2002; Huang et al., 2023; Amin Azad et al., 2020) such attacks, CAPTCHAs remain one of the most cost-effective and widely used defenses.
In addition to protecting web applications, CAPTCHA can also be used to improve the reliability of crowdsourced labeling tasks where there is no objectively correct answer (Alonso et al., 2013), as well as its advanced technique, within-task CAPTCHAs for Human Intelligence Data-Driven Enquiries (HIDDEN) (Alonso et al., 2014) to hold promise as a valuable approach for assessing worker reliability and potentially enhancing label quality.
@@
With the recent surge in multi-modal Large Language Models (MLLMs) (Iong et al., 2024; Yin et al., 2023), artificial intelligence (AI) has gained unprecedented capabilities in handling a wide range of complex tasks, including but not limited to understanding text, multimedia, and generating synthesized conclusions. Furthermore, by granting AI the ability to control computers (Ma et al., 2024), these systems can now take autonomous actions in the digital world. If an AI-powered agent could operate a computer and control the browser like a human, techniques like browser fingerprinting (Eckersley, 2010; Nikiforakis et al., 2013) and canvas fingerprinting (Laperdrix et al., 2019) in preventing bots will face diminishing effectiveness. Further, once an AI bot which integrated by MLLMs learns how to overcome CAPTCHA challenges, the current CAPTCHA systems could become obsolete overnight.
@@
In recent years, AI-driven video generation models have emerged in an exponential manner, with examples such as Stable Video Diffusion (SVD) (Blattmann et al., 2023), Sora [^1] https://openai.com/index/sora/ (Liu et al., 2024), Pika Labs [^2] https://pika.art/, Runway [^3] https://runwayml.com/ (Cui et al., 2024), Stability AI [^4] https://stability.ai/, Kling [^5] https://kling.kuaishou.com/, Mora (Yuan et al., 2024) and others (Ho et al., 2022a, b; Luo et al., 2023; Khachatryan et al., 2023). The videos generated by these models have proliferated on various social media platforms. As model architectures continue to improve and the size of model parameters expands, the length and quality of the generated videos are expected to reach new heights, making it increasingly difficult to distinguish them from real content. In addition to generating videos from prompts and related images, some models are especially capable of video extension (known as video prolongation, and video prediction). Moreover, these models can customize extended scenes, actions, visuals, and objects within the video using text prompts. They can also create visuals that do not exist in the real world, similar to special effects in films.
@@
Motivation. In light of the above observations and based on the user study results from the research on unnature changes in video content (Hénaff et al., 2019) and AI-extended videos (Wang et al., 2025), we argue that the human ability to recognize sudden events, transitions, or anomalies inconsistent with the real world in video content can be harnessed to design a novel web CAPTCHA system. This system would serve as a defense against bot attacks on the internet. The video data used for the CAPTCHA could be generated using AI techniques, specifically by leveraging the extending video capabilities of video generation models.
@@
Approach. We design a novel CAPTCHA mechanism aimed at defending against current and future, more intelligent web bots and crawlers. This mechanism is based on abrupt changes and transitions in video frames and storylines, where users are asked to identify the points of transition. Since collecting large datasets of real videos with significant transitions is costly, and the degree of transition in real videos is limited, we use AI video generation models to guide and extend these transitions. As a result, our CAPTCHA mechanism relies on human perception to identify guided AI-extended video boundaries, and which we name BounTCHA. Based on this design, we develop a prototype whereby, after watching the video, users are asked to drag the progress bar to the point where they believe the transition occurs and submit their response. The collected timing data is used to determine the effective range within which humans can recognize the boundary, which then serves as the basis for distinguishing human users from bots. Lastly, we conduct a detailed security analysis to assess the effectiveness of this defense method.
@@
The following sections review relevant literature, outline video data preparation, detail the prototype and studies on human performance, discuss the in-depth security analysis, and conclude with limitations and future work.
@@ 2.3. Video Understanding
In the realm of video understanding tasks, video foundation models (VFMs) have embarked on promising endeavors in exploring model architectures, as evidenced by references such as (Wang et al., 2023b; Lei et al., 2021; Xu et al., 2024). Additionally, they have made significant progress in learning paradigms, (Li et al., 2023b; Xu et al., 2021; Zellers et al., 2021; Xiao et al., 2024; Yao et al., 2024) serving as examples. As for the latest advancements in video understanding, Tarsier (Wang et al., 2024) emerges as a prominent family of large-scale video-language models. These models are meticulously designed to craft high-quality video descriptions. Notably, the model, along with its code and data, has been made publicly accessible for inference, evaluation, and deployment purposes. As of September 3rd, 2024, extensive evaluation results have demonstrated the remarkable capabilities of Tarsier in general video understanding. It has achieved state-of-the-art (SOTA) performance across six open benchmarks, solidifying its position as a leading approach to understand videos. Consequently, in our current research endeavor, we have decided to utilize this model to participate in the prompt generation of AI-extended videos.
@@ 3. Video Data Preparation (RQ1)
3.1. Video Generation Pipeline
@@ 3.2.2. Prompts for Video Generation Prompts
After obtaining the description and five keywords of the raw video, we incorporate this information into the subsequent video generation prompts, along with the last frame of the raw video. These inputs are then fed into the MLLM to generate a completely new video narrative. This process ensures that the generated prompts significantly differ from the storyline and visuals of the raw video, while still being grounded in its foundational elements. Subsequently, the prompts used for guiding video generation are input into the video generative AI to contribute to the creation of the AI-extended video.
@@ 3.3. Video Quality & Shift Cutting
3.4. Time Cost of Video Production
Figure 5 illustrates the time consumption in various stages of video preparation, such as video understanding for 21.7 seconds, prompt generation for 5.4 seconds, video generation for 521.2 seconds and video compression for 3.1 seconds. Although the average time required to generate a video exceeds 500 seconds, multiple variants can be produced rapidly through shift cutting.
Figure 5. The time cost of the video preparation pipeline. The length of the blocks is not drawn to scale based on the time duration.
@@ 5.1. Experiment
Ethics. The school’s Institutional Review Board (IRB) reviewed and approved this human-subjects research.
Participants. We recruited 186 participants for the experiment by posting flyers on campus and advertising on the social media (comprising 83 participants in-person and 103 participants via ZOOM for online experiments). Participants’ ages ranged from 18 to 48 years ($\mu$ = 26.93, $\sigma$ = 7.57), and all had prior experience with web-based CAPTCHAs. Each participant compensated with \$0.7 USD. To ensure data quality and the accuracy of subsequent research, each participant in each experiment is assigned a supervisor. The supervisor will provide training on the experimental procedures before the experiment begins, address any questions the participant may have during the experiment, and monitor the participant’s attentiveness (i.e., ensuring that the answers submitted are not random).
@@ 5.2. Results
Figure 11. Left: the time bias range and the success rate of the group in the age 18-29. Mid: in the age 30-39. Right: in the age 40-48.
@@ 7. Conclusion
We designed and implemented a novel CAPTCHA, named BounTCHA, which used guided AI-extended videos as the material, to address the growing threat posed by increasingly intelligent AI-powered bots. BounTCHA was designed to differentiate between genuine users and bots based on human perception and identification of time boundaries in video transitions and abrupt changes. We developed a prototype of BounTCHA to conduct experiments and determine the effective range of human perceptual time bias, which serves as a basis for distinguishing between real users and bots. A comprehensive security analysis was then performed on BounTCHA, covering random attacks, database attacks, and multi-modal LLM attacks. The results of the analysis demonstrated that BounTCHA effectively defends against various attack vectors. We envision BounTCHA as a robust shield in web security, safeguarding millions of web applications from AI-powered bot threats in an era where machines are becoming increasingly intelligent.